Three people picked 65 apples altogether. At the first tree they each picked the same number of apples. At the second tree they picked 3 times as many as they picked at the first tree. When they finished at the third tree, the group had 5 times as many apples as they had when they started at the tree. At the fourth tree the group picked just 5 apples.
How many apples did each person pick at the first tree?

Answer by ptaylor(2198)

Let x=number of apples picked by each at the first tree
Then at the first tree, 3x apples were picked
At the second tree, they picked 3*3x=9x apples
When they started at the third tree, the group had 12x apples
when they finished at the third tree, the group had 5*12x=60x apples
So total apples picked =60x+5 and that equals 65
soooo our equation to solve is:
60x+5=65
60x=60
x=1 so they each picked 1 apple at the first tree
CK
first tree=3 apples picked
Second tree=3*3=9 apples
Started at third tree=9+3=12 apples
Finished third tree=5*12=60 apples
60+5=65
65=65
Hope this helps--ptaylor